### v4

@@ 3 Order Finding and Applications
The following problem is in $\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$: On input of a multiplication table of a semigroup $S$, an element $s\in S$, and a unary or binary number $k\in\mathbb{N}$ with $k\leq\left|\mathinner{S}\right|$, compute $s^{k}$ (NB: Definition/Lemma 2.3).
@@
Proof.
If $k$ is given in unary, we first compute its binary representation using a binary search (note that we can write it on the work tape as is uses at most $\left\lceil\mathinner{\log\left|\mathinner{S}\right|}\right\rceil$ bits). We identify the semigroup elements with the natural numbers $10,\dots,\left|\mathinner{S}\right|-$ 1. Now, we compute $s^{k}$ using the standard fast exponentiation algorithm. Note that the multiplication of two semigroup elements can be done in $\mathsf{DTIME}(\log n)$ as we only need to write down two $\log n$ bit numbers on the address tape (if the multiplication table is not padded up to a power of two, this is still in $\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ because we need to multiply two $\log n$ bit numbers to compute the index in the multiplication table).
It is well-known that the fast exponentiation algorithm needs only $O(\log k)$ elementary multiplications and $O(\log k+\log n)$ space; hence, the lemma follows. ∎
Note that Lem. 3.1 together with Lem. 2.5 gives a new proof that the problem of computing a power $s^{k}$ in a semigroup can be done in $\mathsf{FOLL}$. This approach seems easier and more general than the double-barrelled recursive approach in [BKLM01].
@@
For the depth-2 upper bound: we can compute the decision problems $(g,i)\mapsto[g^{i}=1]$, $(g,i)\mapsto[g^{i}\neq 1]$, and $(g,h,i)\mapsto[g^{i}=1\leftrightarrow h^{i}=1]$ in $\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$, which in turn can be decided by quasi-polynomial-size CNFs. Implementing the above strategy, the $\forall^{\log n}$ becomes an AND gate of fan-in $\operatorname{poly}(n)$. That can be merged with the AND gates at the top of the aforementioned CNFs to get a single CNF of quasi-polynomial size. ∎
@@ 3.1 Application to isomorphism testing
Using Cor. 3.2, we can improve the upper bound for isomorphism testing of finite simple groups. Previously, this problem was known to be in $\mathsf{L}$ [Tan13] and FOLL [GL23]. We obtain the following improved bound.
@@
Proof.
As $G$ is a finite simple group, $G$ is determined up to isomorphism by (i) $|G|$, and (ii) the set of orders of elements of $G$: $\text{spec}(G):=\{\operatorname{ord}(g):g\in G\}$ [VGM09]. We may check whether $|G|=|H|$ in $\mathsf{AC}^{0}$. By Cor. 3.2, we may compute and compare $\text{spec}(G)$ and $\text{spec}(H)$ in $\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$. ∎
In light of Cor. 3.2, we obtain an improved bound for testing whether a group $G$ is nilpotent. Testing for nilpotency was previously known to be in $\mathsf{L}\cap\textsf{FOLL}$ [BKLM01].
@@
Proof.
A finite group $G$ is nilpotent if and only if it is the direct product of its Sylow subgroups. As $G$ is given by its multiplication table, we may in $\mathsf{AC^{0}}$ compute the prime factors of $|G|$. Thus, for each prime $p$ dividing $|G|$, we identify the elements $X_{p}:=\{g:\operatorname{ord}(g)\text{ is a power of }p\}$ and then test whether $X_{p}$ forms a group. By Cor. 3.2, we can identify $X_{p}$ in $\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$. Verifying the group axioms is $\mathsf{AC^{0}}$-computable. The result follows. ∎
@@ 3.2 Application to membership testing
A group $G$ has the $\log n$ power basis property (as defined in [BKLM01, Fle22]) if for every subset $X\subseteq G$ every $g\in\langle X\rangle$ can be written as $g=g_{1}^{e_{1}}\cdots g_{m}^{e_{m}}$ with $m\leq\log n$ and suitable $g_{i}\in X$ and $e_{i}\in\mathbb{Z}$.
Observation 3.5.
@@
Proof.
This follows by guessing suitable exponents and using Lem. 3.1 to compute the respective powers. In order to keep the space logarithmically bounded, we do not simply guess $g_{1},\dotsc,g_{m}\in X$—which could take up to $\log^{2}n$ bits—but rather guess the $g_{i}$ sequentially and only store the running product. That is, first guess $g_{1}$ and $e_{1}$ using $2\log n$ bits, and compute $g_{1}^{e_{1}}$. Inductively suppose we have computed $g_{1}^{e_{1}}\dotsb g_{i}^{e_{i}}$. Note that we do not need to store the elements $g_{1},\dotsc,g_{i}$ and exponents $e_{1},\dotsc,e_{i}$, we only keep the product $g_{1}^{e_{1}}\dotsb g_{i}^{e_{i}}$, which takes $\log n$ bits. Then we guess $g_{i+1}$ and $e_{i+1}$, and continue. ∎
This observation allows us to improve several results from [BKLM01, Fle22] from $\mathsf{FOLL}$ to $\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$.
@@
Proof.
Commutative semigroups have the $\log n$ power basis property [Fle22, Lem. 4.2]. ∎
@@
Proof.
Let $X$ be the input set for Membership and $G$ the input group. Theorem 3.5 in [BKLM01] is proved by showing that $\langle X\rangle=X_{d}$ where $X_{0}=X$ and for some suitable constant $C$
@@
Now, as in 3.5, we can guess an element of $X_{i}$ in $\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ given that we have the elements of $X_{i-1}$. As $d$ is a constant, we can guess the elements of $X_{i-1}$ in $\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ by induction. Thus, we can decide membership in $\langle X\rangle=X_{d}$ in $\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$. ∎
For nilpotent groups, we can do even better and show $\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ even without a bound on the nilpotency class—thus, improving considerably over [BKLM01]. This also underlines that the class $\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ is useful not only because it provides a better complexity bound than $\mathsf{FOLL}$, but it also facilitates some proofs.
@@
Proof.
Let $X$ be the input set for Membership and $G$ the input group. Write $n=\left|\mathinner{G}\right|$. Note that the nilpotency class of $G$ is at most $\log n$. Define $C=\{[x_{1},\dots,x_{\ell}]\mid x_{i}\in X,\ell\leq\log n\}$ where $[x_{1},\dots,x_{\ell}]$ is defined inductively by $[x_{1},\dots,x_{\ell}]=[[x_{1},\dots,x_{\ell-1}],x_{\ell}]$ for $\ell\geq 3$.
@@
Let $\Gamma_{m}$ denote the $m$-th term of the lower central series of $\langle X\rangle$ (meaning that $\Gamma_{0}=\langle X\rangle$ and $\Gamma_{m+1}=[\Gamma_{m},\langle X\rangle]$). Then $\Gamma_{m}$ is generated by $C_{m}=\{[x_{1},\dots,x_{k}]\mid x_{i}\in X,k\geq m\}$ (e. g., [CMZ17, Lemma 2.6]). Observe that, although $k$ is unbounded, the terms with $k>\log n$ are trivial because $\log n$ is a bound on the nilpotency class. We obtain the desired set $C^{\prime}$ by first choosing a minimal generating set for the Abelian group $\Gamma_{0}/\Gamma_{1}$, then a minimal generating set of $\Gamma_{1}/\Gamma_{2}$ and so on (meaning that $(c_{1},\dots,c_{m})$ is a so-called polycyclic generating sequence of $\langle X\rangle$). Note that $m$ is bounded by $\log n$ since $\langle c_{i},\dots,c_{m}\rangle$ is a proper subgroup of $\langle c_{i-1},\dots,c_{m}\rangle$ for all $i$.
Next, let us show that we can guess an element of $C^{\prime}$ in $\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$. This is not difficult: we start by guessing $x_{1}$ and set $a=x_{1}$. Then, as long as we guess to continue, we guess $x_{i}$ and compute $a:=[a,x_{i}]$. As $\ell$ (in the definition of $C$) is bounded by $\log n$, we can guess any element of $C^{\prime}$ in this way in polylogarithmic time. Moreover, note that at any point we only need to store a constant number of group elements.
Once we have guessed an element of $C^{\prime}$, we can guess a power of it as in 3.5 and then guess the next element of $C^{\prime}$ and so on. As $m\leq\log n$, this gives an $\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ algorithm. ∎
@@ 4 Abelian Group Isomorphism
Chattopadhyay, Torán, and Wagner [CTW13] established a $\textsf{TC}^{0}(\mathsf{FOLL})$ upper bound on this problem. Grochow & Levet [GL23, Theorem 5] gave a tighter analysis of their algorithm, placing it in the sub-class $\forall^{\log n}\mathsf{MAC^{0}}(\mathsf{FOLL})$. [^15] Grochow & Levet consider $\forall^{\log n}\mathsf{MAC^{0}}\circ\mathsf{FOLL}$, where $\circ$ denotes composition (see [GL23] for a precise formulation). We note that as $\mathsf{AC^{0}}\circ\mathsf{FOLL}=\mathsf{FOLL}=\mathsf{AC^{0}}(\mathsf{FOLL})$, we have that $\forall^{\log n}\mathsf{MAC^{0}}\circ\mathsf{FOLL}=\forall^{\log n}\mathsf{MAC^{0}}(\mathsf{FOLL})$. Thus, Thm. 4.1 improves upon the previous bound of $\forall^{\log n}\mathsf{MAC^{0}}(\mathsf{FOLL})$ obtained by Grochow & Levet. We note that Chattopadhyay, Torán, & Wagner also established an upper bound of $\mathsf{L}$ for this problem, which is incomparable to the result of Grochow & Levet (ibid.). We improve upon both these bounds by (i) showing that $O(\log\log n)$ non-deterministic bits suffice instead of $O(\log n)$ bits, and (ii) using an $\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$ circuit for order finding rather than an $\mathsf{FOLL}$ circuit. We note that while $\forall^{\log n}\mathsf{MAC^{0}}(\mathsf{FOLL})$ is contained in $\mathsf{TC^{0}}(\mathsf{FOLL})$, it is open whether this containment is strict. In contrast, Cor. 4.3 shows that our new bound of $\forall^{\log\log n}\mathsf{MAC}^{0}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$ is a class that is in fact strictly contained in $\mathsf{L}\cap\textsf{TC}^{0}(\mathsf{FOLL})$.
Proof of Thm. 4.1.
Following the strategy of [GL23, Theorem 7.15], we show that non-isomorphism can be decided in the same class but with existentially quantified non-deterministic bits.
We may check in $\mathsf{AC^{0}}$ whether a group is Abelian. So if $H$ is not Abelian, we can decide in $\mathsf{AC^{0}}$ that $G\not\cong H$. So suppose now that $H$ is Abelian. By the Fundamental Theorem of Finite Abelian Groups, $G$ and $H$ are isomorphic if and only if their multiset of orders are the same. In particular, if $G\not\cong H$, then there exists a prime power $p^{e}$ such that there are more elements of order $p^{e}$ in $G$ than in $H$. We first identify the order of each element, which is $\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$-computable by Lem. 3.1.
@@
However, to identify elements of order $p^{e}$, we will need the number $p^{e}$ explicitly, in its full binary expansion. First we show that we can get $p^{e}$ explicitly if we can get $p$ explicitly. Once we have $p$ in its binary expansion, the function $(p,e)\mapsto p^{e}$ is a function of $O(\log n)$ bits, so can be computed by an $\mathsf{AC}^{0}$ circuit of size $2^{O(\log n)}=n^{O(1)}$ (see Thm. 2.2). Thus all that remains is to get $p$ explicitly.
Because all the numbers involved are only $O(\log n)$ bits, any arithmetic function of these numbers can be computed in $\mathsf{AC}^{0}$, in particular, testing if an $O(\log n)$-bit number is prime, testing if one $O(\log n)$-bit number divides another. So, in parallel, for all numbers $x=2,3,\dotsc,n/2$, an $\mathsf{AC}^{0}$ checks which ones are prime and divide $n$. Now, consider the surviving primes as a list of length $O(\log n)$, consisting of numbers each of $O(\log n)$ bits. Now for each pair of primes $p_{j},p_{h}$, we define an indicator $X(j,h)=1\iff p_{j}>p_{h}$. As $p_{j},p_{h}$ are representable using $O(\log n)$ bits, we may compute $X(j,\ell)$ in $\textsf{AC}^{0}$. Now as the number of primes $\ell\leq\log_{2}(n)$, we may in $\mathsf{AC^{0}}$ find a prime $j$ with:
@@
Theorem 4.2 (Barrington & Straubing [BS94, Thm. 7]).
Let $k>1$. Any $\mathsf{TC}$ circuit family of constant depth, size $2^{n^{o(1)}}$, and with at most $n^{o(1)}$ Majority gates cannot compute the $\textsc{MOD}_{k}$ function.
@@
Proof.
Let $L\in Q^{o(\log n)}\mathsf{MAC}^{0}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$. By Fact 2.4, $\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))\subseteq\mathsf{quasiAC^{0}}$, so $\mathsf{MAC}^{0}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))\subseteq\mathsf{quasiMAC^{0}}$, that is, quasi-polynomial size circuits of constant depth with a single Majority gate at the output. Thus $L\in Q^{o(\log n)}\mathsf{quasiMAC}^{0}$.
@@
Now, to compute $L$, the quantifiers $Q^{o(\log n)}$ can be replaced by a constant-depth circuit (whose depth is equal to the number of quantifier alternations), and whose total size is $2^{o(\log n)}=n^{o(1)}$, where at each leaf of this circuit, we put the corresponding circuit $C_{y}$. The resulting circuit is a $\mathsf{quasiTC}^{0}$ circuit with only $n^{o(1)}$ Majority gates, hence by Theorem 4.2, cannot compute Parity. ∎

### v5

@@ 3 Order Finding and Applications
The following function is in $\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ (NB: Definition/Lemma 2.5): On input of a multiplication table of a semigroup $S$, an element $s\in S$, and a unary or binary number $k\in\mathbb{N}$ with $k\leq\left|\mathinner{S}\right|$, compute $s^{k}$.
@@
Proof 3.2.
If $k$ is given in unary, we first compute its binary representation using a binary search (note that we can write it on the work tape as is uses at most $\left\lceil\mathinner{\log\left|\mathinner{S}\right|}\right\rceil$ bits). We identify the semigroup elements with the natural numbers $0,\dots,\left|\mathinner{S}\right|-1$. Now, we compute $s^{k}$ using the standard fast exponentiation algorithm. Note that the multiplication of two semigroup elements can be done in $\mathsf{DTIME}(\log n)$ as we only need to write down two $\log n$ bit numbers on the address tape (if the multiplication table is not padded up to a power of two, this is still in $\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ because we need to multiply two $\log n$ bit numbers to compute the index in the multiplication table).
It is well-known that the fast exponentiation algorithm needs only $O(\log k)$ elementary multiplications and $O(\log k+\log n)$ space; hence, the lemma follows.
Note that Lem. 3.1 together with Lem. 2.8 gives a new proof that the problem of computing a power $s^{k}$ in a semigroup can be done in $\mathsf{FOLL}$. This approach seems easier and more general than the double-barrelled recursive approach in [10].
@@
For the depth-2 upper bound: we can compute the decision problems $(g,i)\mapsto[g^{i}=1]$, $(g,i)\mapsto[g^{i}\neq 1]$, and $(g,h,i)\mapsto[g^{i}=1\leftrightarrow h^{i}=1]$ in $\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$, which in turn can be decided by quasi-polynomial-size CNFs. Implementing the above strategy, the $\forall^{\log n}$ becomes an AND gate of fan-in $\operatorname{poly}(n)$. That can be merged with the AND gates at the top of the aforementioned CNFs to get a single CNF of quasi-polynomial size.
@@ 3.1 Application to isomorphism testing
Using Cor. 3.3, we can improve the upper bound for isomorphism testing of finite simple groups. Previously, this problem was known to be in $\mathsf{L}$ [53] and $\mathsf{FOLL}$ [26]. We obtain the following improved bound.
@@
Proof 3.6.
As $G$ is a finite simple group, $G$ is determined up to isomorphism by (i) $|G|$, and (ii) the set of orders of elements of $G$: $\text{spec}(G):=\{\operatorname{ord}(g):g\in G\}$ [56]. We may check whether $|G|=|H|$ in $\mathsf{AC}^{0}$. By Cor. 3.3, we may compute and compare $\text{spec}(G)$ and $\text{spec}(H)$ in $\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$.
In light of Cor. 3.3, we obtain an improved bound for testing whether a group $G$ is nilpotent. Testing for nilpotency was previously known to be in $\mathsf{L}\cap\mathsf{FOLL}$ [10].
@@
Proof 3.8.
A finite group $G$ is nilpotent if and only if it is the direct product of its Sylow subgroups. As $G$ is given by its multiplication table, we may in $\mathsf{AC^{0}}$ compute the prime factors of $|G|$ (Lemma 2.3). Thus, for each prime $p$ dividing $|G|$, we identify the elements $X_{p}:=\{g:\operatorname{ord}(g)\text{ is a power of }p\}$ and then test whether $X_{p}$ forms a group. By Cor. 3.3, we can identify $X_{p}$ in $\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$. Verifying the group axioms is $\mathsf{AC^{0}}$-computable. The result follows.
@@ 3.2 Application to membership testing
A group $G$ has the $\log n$ power basis property (as defined in [10, 25]) if for every subset $X\subseteq G$ every $g\in\langle X\rangle$ can be written as $g=g_{1}^{e_{1}}\cdots g_{m}^{e_{m}}$ with $m\leq\log n$ and suitable $g_{i}\in X$ and $e_{i}\in\mathbb{Z}$.
{observation}
@@
Proof 3.9.
This follows by guessing suitable exponents and using Lem. 3.1 to compute the respective powers. In order to keep the space logarithmically bounded, we do not simply guess $g_{1},\dotsc,g_{m}\in X$—which could take up to $\log^{2}n$ bits to store—but rather guess the $g_{i}$ sequentially and only store the running product. That is, first guess $g_{1}$ and $e_{1}$ using $2\log n$ bits, and compute $g_{1}^{e_{1}}$. Inductively suppose we have computed $g_{1}^{e_{1}}\dotsb g_{i}^{e_{i}}$. Note that we do not need to store the elements $g_{1},\dotsc,g_{i}$ and exponents $e_{1},\dotsc,e_{i}$, we only keep the product $g_{1}^{e_{1}}\dotsb g_{i}^{e_{i}}$, which takes $\log n$ bits. Then we guess $g_{i+1}$ and $e_{i+1}$, and continue.
This observation allows us to improve several results from [10, 25] from $\mathsf{FOLL}$ to
$\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$.
@@
Proof 3.11.
Commutative semigroups have the $\log n$ power basis property [25, Lem. 4.2].
@@
Proof 3.13.
Let $X$ be the input set for Membership and $G$ the input group. Theorem 3.5 in [10] is proved by showing that $\langle X\rangle=X_{d}$ where $X_{0}=X$ and for some suitable constant $C$
@@
Now, as in Section 3.2, we can guess an element of $X_{i}$ in $\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ given that we have the elements of $X_{i-1}$. As $d$ is a constant, we can guess the elements of $X_{i-1}$ in $\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ by induction. Thus, we can decide membership in $\langle X\rangle=X_{d}$ in $\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$.
For nilpotent groups, we can do even better and show $\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ even without a bound on the nilpotency class—thus, improving considerably over [10]. This also underlines that the class $\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ is useful not only because it provides a better complexity bound than $\mathsf{FOLL}$, but it also facilitates some proofs.
@@
Proof 3.15.
Let $X$ be the input set for Membership and $G$ the input group. Write $n=\left|\mathinner{G}\right|$. Note that the nilpotency class of $G$ is at most $\log n$. Define $C=\{[x_{1},\dots,x_{\ell}]\mid x_{i}\in X,\ell\leq\log n\}$ where $x_{1},\dots,x_{\ell}]$ is defined inductively by $x_{1},\dots,x_{\ell}]=[[x_{1},\dots,x_{\ell-1}],x_{\ell}]$ for $\ell\geq 3$. When $\ell=1$, define $x_{1},\ldots,x_{\ell}]:=x_{1}$; and when $\ell=2$, $x_{1},\ldots,x_{\ell}]:=[x_{1},x_{2}]$.
@@
Let $\Gamma_{m}$ denote the $m$-th term of the lower central series of $\langle X\rangle$ (meaning that $\Gamma_{0}=\langle X\rangle$ and $\Gamma_{m+1}=[\Gamma_{m},\langle X\rangle]$). Then $\Gamma_{m}$ is generated by $C_{m}=\{[x_{1},\dots,x_{k}]\mid x_{i}\in X,k\geq m\}$ (e. g., [18, Lemma 2.6]). Observe that, although $k$ is unbounded, the terms with $k>\log n$ are trivial because $\log n$ is a bound on the nilpotency class. We obtain the desired set $C^{\prime}$ by first choosing a minimal generating set for the Abelian group $\Gamma_{0}/\Gamma_{1}$, then a minimal generating set of $\Gamma_{1}/\Gamma_{2}$ and so on (meaning that $(c_{1},\dots,c_{m})$ is a so-called polycyclic generating sequence of $\langle X\rangle$– see [31, Chapter 8]). Note that $m$ is bounded by $\log n$ since $\langle c_{i},\dots,c_{m}\rangle$ is a proper subgroup of $\langle c_{i-1},\dots,c_{m}\rangle$ for all $i$.
Next, let us show that in $\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ we can non-deterministically guess elements such that (1) all guessed elements are in $\langle X\rangle$ and (2) every power-product of elements of $C^{\prime}$ occurs on at least one non-deterministic branch.
This is not difficult: we start by guessing $x_{1}$ and set $a=x_{1}$. Then, guess whether or not to continue; as long as we guess to continue, we guess $i$ and compute $a:=[a,x_{i}]$. As $\ell$ (in the definition of $C$) is bounded by $\log n$, we can guess any element of $C^{\prime}$ in this way in polylogarithmic time. Moreover, note that at any point we only need to store a constant number of group elements. By construction, all such guessed elements are visibly in $\langle X\rangle$.
Once we have guessed an element of $C^{\prime}$, we can guess a power of it as in Section 3.2 and then guess the next element of $C^{\prime}$ and so on. As $m\leq\log n$, this gives an $\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ algorithm.
@@ 4 Abelian Group Isomorphism
Chattopadhyay, Torán, and Wagner [17] established a $\mathsf{TC}^{0}(\mathsf{FOLL})$ upper bound on this problem. Grochow & Levet [26, Theorem 5] gave a tighter analysis of their algorithm, placing it in the sub-class $\forall^{\log n}\mathsf{MAC^{0}}(\mathsf{FOLL})$. [^15] Grochow & Levet consider $\forall^{\log n}\mathsf{MAC^{0}}\circ\mathsf{FOLL}$, where $\circ$ denotes composition (see [26] for a precise formulation). We note that as $\mathsf{AC^{0}}\circ\mathsf{FOLL}=\mathsf{FOLL}=\mathsf{AC^{0}}(\mathsf{FOLL})$, we have that $\forall^{\log n}\mathsf{MAC^{0}}\circ\mathsf{FOLL}=\forall^{\log n}\mathsf{MAC^{0}}(\mathsf{FOLL})$. Thus, Thm. 4.1 improves upon the previous bound of $\forall^{\log n}\mathsf{MAC^{0}}(\mathsf{FOLL})$ obtained by Grochow & Levet. We note that Chattopadhyay, Torán, & Wagner also established an upper bound of $\mathsf{L}$ for this problem, which is incomparable to the result of Grochow & Levet (ibid.). We improve upon both these bounds by (i) showing that $O(\log\log n)$ non-deterministic bits suffice instead of $O(\log n)$ bits, and (ii) using an $\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$ circuit for order finding rather than an $\mathsf{FOLL}$ circuit. We note that while $\forall^{\log n}\mathsf{MAC^{0}}(\mathsf{FOLL})$ is contained in $\mathsf{TC^{0}}(\mathsf{FOLL})$, it is open whether this containment is strict. In contrast, Cor. 4.4 shows that our new bound of $\forall^{\log\log n}\mathsf{MAC}^{0}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$ is a class that is in fact strictly contained in $\mathsf{L}\cap\mathsf{TC}^{0}(\mathsf{FOLL})$.
Jeřábek also previously established bounds of $\Sigma_{2}$-$\mathsf{TIME}(\log^{2}n)$ for isomorphism testing of Abelian groups, which can be simulated by depth-$3$ $\mathsf{quasiAC^{0}}$-circuits of size $n^{O(\log n)}$ [34]; Jeřábek’s result and our Theorem 4.1 are incomparable.
Proof 4.2 (Proof of Thm. 4.1).
Following the strategy of [26, Theorem 7.15], we show that being non-isomorphic can be decided in the same class but with existentially quantified non-deterministic bits.
We may check in $\mathsf{AC^{0}}$ whether a group is Abelian. So if $H$ is not Abelian, we can decide in $\mathsf{AC^{0}}$ that $G\not\cong H$. So suppose now that $H$ is Abelian. By the Fundamental Theorem of Finite Abelian Groups, $G$ and $H$ are isomorphic if and only if their multisets of orders are the same. In particular, if $G\not\cong H$, then there exists a prime power $p^{e}$ such that there are more elements of order $p^{e}$ in $G$ than in $H$. We first identify the order of each element, which is $\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$-computable by Lem. 3.1.
@@
However, to identify elements of order $p^{e}$, we will need the number $p^{e}$ explicitly, in its full binary expansion. First we show that we can get $p^{e}$ explicitly if we can get $p$ explicitly. Once we have $p$ in its binary expansion, the function $(p,e)\mapsto p^{e}$ can be computed in $\mathsf{AC}^{0}$, by Thm. 2.2. Thus all that remains is to get $p$ explicitly.
By Lemma 2.3 we can compute in $\mathsf{AC^{0}}$ the prime factors of $|G|$. Consider this list of primes as a list of length $O(\log n)$, consisting of numbers each of $O(\log n)$ bits. Now for each pair of primes $p_{j},p_{h}$, we define an indicator $X(j,h)=1\iff p_{j}>p_{h}$. As $p_{j},p_{h}$ are representable using $O(\log n)$ bits, we may compute $X(j,\ell)$ in $\mathsf{AC}^{0}$. Now as the number of primes $\ell\leq\log_{2}(n)$, we may in $\mathsf{AC^{0}}$ find a prime $j$ with:
@@
Theorem 4.3 (Barrington & Straubing [12, Thm. 7]).
Let $k>1$. Any $\mathsf{TC}$ circuit family of constant depth, size $2^{n^{o(1)}}$, and with at most $n^{o(1)}$ $\mathsf{Majority}$ gates cannot compute the $\textsc{Mod}_{k}$ function.
@@
Proof 4.5.
Let $L\in Q^{o(\log n)}\mathsf{MAC}^{0}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$. Since by Fact 1 we know that $\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))\subseteq\mathsf{quasiAC^{0}}$, we have $\mathsf{MAC}^{0}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))\subseteq\mathsf{quasiMAC^{0}}$, that is, quasi-polynomial size circuits of constant depth with a single $\mathsf{Majority}$ gate at the output. Thus $L\in Q^{o(\log n)}\mathsf{quasiMAC}^{0}$.
@@
Now, to compute $L$, the quantifiers $Q^{o(\log n)}$ can be replaced by a constant-depth circuit (whose depth is equal to the number of quantifier alternations), and whose total size is $2^{o(\log n)}=n^{o(1)}$, where at each leaf of this circuit, we put the corresponding circuit $C_{y}$. The resulting circuit is a $\mathsf{quasiTC}^{0}$ circuit with only $n^{o(1)}$ $\mathsf{Majority}$ gates, hence by Theorem 4.3, cannot compute Parity.